Two trains left a station at the same time and traveled in opposite directions. One train traveled for 144 minutes at a rate that was 5 kilometers per hour faster than the speed of the first train. The faster train went twice as far as the slower train. How far did each train travel?
Answer by mananth(16949)

let the speed of train A be x km/h
Train B traveled at x+5 km/h

Train B traveled 144 minutes= 2.4 hours (144/60)
Distance traveled by train B = (x+5)*2.4
Distance traveled by Train A = 2.4x
Distance traveled by train B = twice distance traveled by train A
2.4(x+5)= 2*2.4x
2.4x+12=4.8x
4.8x-2.4x=12
2.4x=12
x=12/2.4
x=5
speed of Train A = 5km/h---------- distance traveled = 2.4*5 12 km
speed of train B = 10km/h----------distance traveled = 10*2.4 = 24 km